OpenStudy (anonymous):

Which statement best describes a chord of a circle? It is a line that touches only one point on the circle. It is a segment drawn from a point on the circle to the center of the circle. It is a segment that connects two distinct points on a circle. It is a line drawn always through the center of the circle.

OpenStudy (anonymous):

It is a segment that connects two distinct points on a circle.

OpenStudy (anonymous):

What about this question? A circle has its center at (-2, -3) and a radius of 3 units. What is the equation of the circle? (x + 2)2 + (y + 3)2= 9 (x + 2)2 + (y - 3)2= 9 (x - 2)2 + (y - 3)2 = 3 (x + 2)2 + (y + 3)2 = 3

OpenStudy (anonymous):

@Calcmathlete

OpenStudy (anonymous):

(x + 2)2 + (y + 3)2= 9